Jackson Company applies overhead to products using direct labor hours
as the activity level. During 2019, Jackson Company had the following
estimated costs:

Direct materials .................... $258,000
Direct labor ........................  255,000
Advertising .........................   35,000
Rent on factory building ............   36,000
Depreciation .................... ...   60,000
Indirect materials ..................   32,000
Sales commissions ...................   50,000
Production supervisor's salary ......   40,000
Insurance on factory equipment ......   18,000
CEO's salary ........................   95,000

The following additional information is available:

1. 30% of the depreciation relates to equipment in the administrative
   offices while 70% of the depreciation relates to equipment in the
   factory.

2. Jackson Company estimated 15,000 direct labor hours would be worked
   during 2019.

During 2019, Jackson Company began work on three jobs. Information
relating to these three jobs appears below:

                                  Job #359      Job #360      Job #361
direct materials ..............    $98,000       $75,000       $91,000
direct labor cost .............    $98,600       $79,900       $86,700
direct labor hours ............      5,800         4,700         5,100

By the end of 2019, job #359 and job #361 had been completed. Job #360
was not completed by the end of 2019. Additionally, by the end of 2019,
job #361 had been sold while job #359 was not sold. Jackson Company had
total actual overhead cost of $169,000 during 2019.

Part 1) Calculate Jackson Company's pre-determined overhead rate (PDR) for 2019 based on direct labor hours.

Part 2) Calculate the amount of finished goods inventory reported in Jackson Company's December 31, 2019 balance sheet.

Part 3) Calculate the cost of goods sold reported by Jackson Company for 2019 after the overhead variance has been closed.
Answer #1

1) Solution: $11.20

Explanation: Estimated overhead cost = [Rent on factory building + Depreciation at 70% + Indirect materials + Insurance on factory equipment + Production supervisor’s salary]

= [36,000 + (60,000 * 70%) + 20,000 + 40,000+18,000 ]

= $156,000

Direct labor are paid at: [$255,000 / 15,000 hours] = $17

Pre-determined OH rate = [$156,000 / 15,000] = $10.40 per direct labor hour

 

2) Solution: $256,920

Explanation: Finished goods indicates the completed units that have not yet been sold.

Job #359 Cost = [98,000 + 98,600 + (5,800 * $10.40)] = $256,920

 

3) Solution: $237,500

Working: Total applied OH = [5,800 + 4,700 + 5,100] * $10.40 = $162,240

OH underapplied = 169,000 - 162,240 = 6,760

Unadjusted COGS indicates the cost of jobs that were sold.

Job #361 Cost = [91,000 + 86,700 + (5,100 * $10.40)] = $230740

Adjusted cost of goods sold = [Unadjusted COGS + Under-applied overhead] = [$230740 + 6,760] = $237,500